OpenStudy (anonymous):

Find a function whose graph is a parabola with vertex (7, 8) and that passes through the point (6, −9).

OpenStudy (anonymous):

The equation in vertex form is: \(\large y=a(x-h)^2+k \) where (h, k) is your vertex. Since (7, 8) is given as the vertex, then the equation looks like this: \(\large y=a(x-7)^2+8 \) But you're also given that the parabola passes through (6, -9) so: \(\large -9=a(6-7)^2+8 \) Solve for "a" then put that back into the equation.

OpenStudy (anonymous):

Wow. That was a lot easier that I was making it. Thank you very much!

hero (hero):

I'm pretty sure my instructor showed me the hard way of doing it. You're lucky AdriRules